OpenStudy (allymariex3):

Simplify the given expression. √ -6 * √ -18 A. 6√ 3 B. -6√ 3 C. 6i√ 3 D. -6i√ 3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

is not a is b

Parth (parthkohli):

Hint:\[\rm = i\sqrt 6 \times i\sqrt{18}\]Continue from this point, and never go with the sqrt a * sqrt b = sqrt ab approach.

OpenStudy (anonymous):

b